Verify that the given family of functions solves the differential equation.
i. dy/dt = (1 - 2t)y² , y = 1/( C - t + t²)
0
ⁱⁱ· dy/dt = y²sin(t) ,y = 1/(C + cos(t))
1
Expert's answer
2020-06-28T08:15:03-0400

(i)

Y= 1/c-t+t2

dy/dt = (1-2t) y2

dy/dt =-1/(c-t+t2)2 ×(-1+2t)

= (1-2t)×1/(c-t+t2)2

dy/dt =(1-2t)×y2 => answer


(ii)

y = 1/(c+cost),

dy/dt= -1/(c+cost)2×(-sint),

dy/dt =(sint)y2 => answer